OpenStudy (boxman61):

been struggling for hours.. Solve: log(2)3+log(2)x=log(2)(x+7) please help me solve

OpenStudy (boxman61):

\[\log_{2}3+\log_{2}x=\log_{2}(X+7) \]

OpenStudy (anonymous):

log(A*B)=logA + logB now u cn solve?

OpenStudy (boxman61):

well i've done \[\log_{2}3+\log_{2}x=\log_{2}(x+7) \]\[3x=(x+7)\]\[2x=7\]\[x=7/2\] but is that the correct math?

OpenStudy (anonymous):

yeahh its absolutely correct as base of log is same on both sides...

OpenStudy (boxman61):

so the answer really is x=7/2, thank you

OpenStudy (anonymous):

yeahhh absolutely correct...